A lorry was travelling at a constant 50mph. I he travelled for 1 1/2 hours, how far will he have travelled

Answer by richwmiller(17219)

Glad you realized it is a word problem but you couldn't tell it involves travel and distance?
in the first hour the lorry went 50 miles that leaves half an hour
in the half an hour he went 50/2=25
so altogether he went 50+25=75
50*1.5=75
r*t=d
rate *time=distance